OpenStudy (anonymous):

what is the value of e^-2

OpenStudy (anonymous):

\[\frac {1} {e^2} = \frac {1} {2.71^2} \]

OpenStudy (anonymous):

approximately equal

OpenStudy (anonymous):

0.135

OpenStudy (anonymous):

well i am looking for the volume so i have (1/2 e^-2 + 1/2 E^0)

OpenStudy (anonymous):

i know the second one goes to 1/2

OpenStudy (anonymous):

what about the firstn one?